Sara just has just rented her first apartment starting December 1 before beginning college in January. The apartment has washer and dryer hook-ups so sara wanted to buy the appliances to avoid trips to the laundomat. The saturday newspaper had an advertisement for a local appliance store offering 90 days same as cash financing sara asked how the financing worked and learned that she could pay for the washer and dryer anytime during the first 90 days for the purchase price plus sales tax,If she waited longer she would have to pay the purchse price, plus sales tax plus 26.8% annual simple interest for the first 90 days plus 2% simple interest per month on the unpaid balance after 90 days, with the minimum payments of 50.00 per month after the 90 day period. Together the washer and dryer cost 699plus the 8.25% sale tax. Sara knew that her tax refund from the IRS would be 1,000 so she bought the washer and dryer confident that she could pay off the balance within the 90 days.
1. If Sara pays off the balance within 90 days how much will she pay?
2. If sara bought the washer and dryer on December 15 using the exact interest, what is her deadline for paying no interest in a non leap year. in a leap year? Is finance company likely to use exact or ordinary interest and why?
3. If sara irs refund does not come until April 1 what is her pay off amount ( Assume ordinary interest and in a non leap year?
4. How much did it cost her to pay off this loan 17 days late ? What annual simple interest rate does this amount to?

Answer by rfer(16322)

1) 699*1.0825=756.67
2) March 14, March 13 will use ordinary for ease of calculations
3) 756.67*1.067=807.37+(.01133*756.67)=815.94
I converted the 2% per month to 17 days of interest or .01133 for that part of the month.
4) $59.27, 26.9% annual rate or 7.8% for the period.